At an output level of 18,500 units, you have calculated that the degree of operating leverage is 3.20. The operating cash flow is $48,000 in this case. Ignoring the effect of taxes, what are fixed costs? (Do not round intermediate calculations. Round your answer to the nearest whole number, e.g., 32.) Fixed costs $ What will the operating cash flow be if output rises to 20,500 units? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Operating cash flow $ What will the operating cash flow be if output falls to 17,000 units? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Operating cash flow $
Answer #1

At an output level of 18,500 units, the degree of operating leverage is 3.20. The operating cash flow is $48,000

Degree of operating leverage (DOL)

= (Sales – Variable cost)/ (Sales – Variable cost – Fixed cost)

= (operating cash flow +fixed cost)/ (operating cash flow)

Fixed cost = (DOL -1) * operating cash flow

= (3.20 -1) * $48,000 = $105,600

Therefore fixed costs is $105,600

Contribution margin per unit = (operating cash flow + fixed cost)/ output level

= ($48,000 + $105,600) / 18,500

= $153,600/ 18,500 = $8.3027

What will the operating cash flow be if output rises to 20,500 units?

At 20,500 units

Operating cash flow = Contribution margin per unit *20,500 - fixed cost

= $8.3027 *20,500 units - $105,600

= $170,205.41 - $105,600

= $64,605.41

Operating cash flow is $64,605.41

What will the operating cash flow be if output falls to 17,000 units?

At 17,000 units

Operating cash flow = $8.3027 *17,000 units - fixed cost

= $141,145.95 - $105,600

= $35,545.95

Operating cash flow is $35,545.95